OpenStudy (anonymous):

if log55 = 1.74 what is the value of log(1000)^55 ?

OpenStudy (anonymous):

log 1000^55 = (log 10^3)^55 = 1/3 log 55 to the base 10 = 1/3 x 1.74 = 0.58

OpenStudy (anonymous):

okay thanks!!